OpenStudy (anonymous):

If you draw a card at random from a deck of playing cards what is the probability it is a heart, spade, club or diamond?

OpenStudy (anonymous):

@jim_thompson5910 can u help me?

jimthompson5910 (jim_thompson5910):

The probability is 1 or 100% because there are only four suits and they are hearts, spades, clubs or diamonds. So it's guaranteed you're going to draw one of the four options.

OpenStudy (anonymous):

okay and can u help me out with a couple more. I need help plese,and your good at explaining.

jimthompson5910 (jim_thompson5910):

sure whats up

OpenStudy (anonymous):

If you dial a seven digit number at random on a phone, what is the probability you dial your next door neighbor's main phone? for this is the answer 1/5040

jimthompson5910 (jim_thompson5910):

Assuming all digits are likely to be dialed at any one time, there are 10^7 = 10,000,000 possible phone numbers to dial But the neighbor only has one phone number (assuming s/he has only one land-line or cell). So the probability is 1/10000000

jimthompson5910 (jim_thompson5910):

Your answer is assuming that you can't re-use any digits but numbers like 555-1212 are perfectly possible.

OpenStudy (anonymous):

oh okay and for thjis question : A shootout in hockey consists of Teams A and B taking alternate shots on goal. Suppose that the rules state that the first team to score wins. Team A has a probability of 0.39 of scoring with any one shot. Team B has a probability of 0.39 of scoring with any one shot. If team A shoots first, what is the probability, to 2 decimal places, of Team B winning on its first shot? is the answer 0.24

jimthompson5910 (jim_thompson5910):

Team A shoots first So the probability of team A winning on the first shot is given as 0.39 In order for B to win on the first shot, A must NOT win on the first shot (otherwise, the game is over) So the probability of A NOT winning on the first shot is 1 - 0.39 = 0.61 This then means that... P(B winning on first shot) = P(A NOT winning on first shot AND B winning on first shot) P(B winning on first shot) = P(A NOT winning on first shot)*P(B winning on first shot) P(B winning on first shot) = 0.61*0.39 P(B winning on first shot) = 0.2379 P(B winning on first shot) = 0.24 So you are correct

OpenStudy (anonymous):

okay and what about this question: City Council consists of 8 men and 5 women. If 6 representatives are chosen at random to form an environmental sub-committee, what is the probability that Mayor Mr. Bill Tompson and 3 women are chosen, correct to the nearest thousandth?

jimthompson5910 (jim_thompson5910):

There are 8 men + 5 women = 13 people total So there are 13 C 6 = 1716 different ways to form a sub-committee (with no restrictions) Since the mayor must be on the sub-committee, there are now 6-1 = 5 places left to fill. There are also 13 - 1 = 12 people left to choose from. There are 5 women, and 3 women must be on the committee, so there are 5 C 3 = 10 different ways to add on 3 women. The remaining 2 spots left and these spots must be filled by men (since there must be exactly 3 women). So there are 7 C 2 = 21 ways to do this So overall there are 1*10*21 = 210 different ways to form the committee that has exactly 3 women and the mayor (and 2 other guys) So the probability is 210/1716 = 35/286 = 0.1223776

OpenStudy (anonymous):

From a group of 13 student representatives, 5 will be chosen to work on a committee. What is the probability that Sally is chosen, correct to the nearest thousandth? is the answer 1/1287

OpenStudy (anonymous):

so the answer to the nearest thousandth would be 0.001

jimthompson5910 (jim_thompson5910):

There are 13 C 5 = 1287 different ways to form this committee Sally must be chosen, so there are 5-1 = 4 slots left. There are 13-1 = 12 students left to choose from. So there are 12 C 4 = 495 ways to choose a committee where Sally is on it. So the probability is 495/1287 = 5/13 = 0.3846

OpenStudy (anonymous):

oh okay i realized my mistake. and 2 last question if you dont mind: At a family reunion, door prizes are to be given out. At one table in the community hall, 6 children, 3 teenagers, 5 adults, and 6 seniors are seated. The 3 winning tickets are held by 3 different people at this table. is the answer for this 0.045

OpenStudy (anonymous):

oh never mind i got it thanks for the help, ill probably be asking u more questions if your sticking around. lol

jimthompson5910 (jim_thompson5910):

It'll depend really, so try to get as much done and ask the ones that are really bugging you.

OpenStudy (anonymous):

hey can u help?

OpenStudy (anonymous):

A basketball player has scored on 60% of her free throws throughout her career. Given this information, the probability that she will score on exactly 9 of her next 14 free throws, correct to the nearest hundredth, is

jimthompson5910 (jim_thompson5910):

Use the binomial distribution probability formula

jimthompson5910 (jim_thompson5910):

P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 9) = (14 C 9)*(0.6)^(9)*(1-0.6)^(14-9) P(X = 9) = (14 C 9)*(0.6)^(9)*(0.4)^(14-9) P(X = 9) = (2002)*(0.6)^(9)*(0.4)^5 P(X = 9) = (2002)*(0.010077696)*(0.01024) P(X = 9) = 0.20659760529408 So the probability that she will score on exactly 9 of her next 14 free throws is roughly 0.20659760529408

OpenStudy (anonymous):

okay thts wht i gowht i got too how abt this one: An examination consists of 50 multiple choice questions each with four possible answers. What is the probability of guessing exactly 10 correct answers?

jimthompson5910 (jim_thompson5910):

Again use the binomial distribution probability formula P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 10) = (50 C 10)*(0.25)^(10)*(1-0.25)^(50-10) P(X = 10) = (50 C 10)*(0.25)^(10)*(0.75)^(50-10) P(X = 10) = (10272278170)*(0.25)^(10)*(0.75)^40 P(X = 10) = (10272278170)*(0.00000095367)*(0.000010056585) P(X = 10) = 0.09851796245376 So the probability of guessing exactly 10 correct answers is 0.09851796245376

OpenStudy (anonymous):

am i bugging u buy asking all these questions,cause i can ask someone else? cause i have 3 more questions to ask.

jimthompson5910 (jim_thompson5910):

you can ask me, i'm just going to be slow to respond because I'm answering other questions

jimthompson5910 (jim_thompson5910):

so it's up to you

OpenStudy (anonymous):

okay ill ask u then: A binomial experiment consists of 6 trials with a probability of success on each trial of 0.2. Calculate the probability of obtaining at least 3 successes.

OpenStudy (anonymous):

is the answer 0.099

OpenStudy (anonymous):

Approximately 3% of peanuts in a certain shipment are spoiled. If you randomly select 24 peanuts, what is the probability that at least one peanut is spoiled? and this one

jimthompson5910 (jim_thompson5910):

one question at a time please

jimthompson5910 (jim_thompson5910):

it's hard to reference multiple questions

OpenStudy (anonymous):

oh sorry about that.

jimthompson5910 (jim_thompson5910):

no worries, this is the answer to the first one P(X >= 3) = P(X = 4) + P(X = 5) + P(X = 6) So we need the following P(X = 4), P(X = 5), and P(X = 6) P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 4) = (6 C 4)*(0.2)^(4)*(1-0.2)^(6-4) P(X = 4) = (6 C 4)*(0.2)^(4)*(0.8)^(6-4) P(X = 4) = (15)*(0.2)^(4)*(0.8)^2 P(X = 4) = (15)*(0.0016)*(0.64) P(X = 4) = 0.01536 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 5) = (6 C 5)*(0.2)^(5)*(1-0.2)^(6-5) P(X = 5) = (6 C 5)*(0.2)^(5)*(0.8)^(6-5) P(X = 5) = (6)*(0.2)^(5)*(0.8)^1 P(X = 5) = (6)*(0.00032)*(0.8) P(X = 5) = 0.001536 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 6) = (6 C 6)*(0.2)^(6)*(1-0.2)^(6-6) P(X = 6) = (6 C 6)*(0.2)^(6)*(0.8)^(6-6) P(X = 6) = (1)*(0.2)^(6)*(0.8)^0 P(X = 6) = (1)*(0.000064)*(1) P(X = 6) = 0.000064 -------------------------------------------- P(X >= 3) = P(X = 4) + P(X = 5) + P(X = 6) P(X >= 3) = 0.01536 + 0.001536 + 0.000064 P(X >= 3) = 0.01696 So the probability of getting at least 3 successes is 0.01696

jimthompson5910 (jim_thompson5910):

Well I guess I could do "Question: ...., then Answer:...", but it's still a good idea to do this one at a time. Question: Approximately 3% of peanuts in a certain shipment are spoiled. If you randomly select 24 peanuts, what is the probability that at least one peanut is spoiled? ---------------------------------------------------------------------- Answer: P(At least one spoiled) = 1 - P(None spoiled) P(At least one spoiled) = 1 - (1-0.03)^24 P(At least one spoiled) = 1 - (0.97)^24 P(At least one spoiled) = 1 - 0.48141722191172 P(At least one spoiled) = 0.51858277808828

OpenStudy (anonymous):

okay thankyou so much jim, i really appreciate it!

jimthompson5910 (jim_thompson5910):

you're welcome